### v2

@@ 4.1 An algorithm to find elements of the rectified exclusion core
We present an algorithm called generalized top trading cycle (GTTC), and prove that all allocations it finds belong to the rectified exclusion core. This implies that the rectified exclusion core is nonempty.
Under strict preferences, TTC is an individually rational, Pareto efficient, and strategy-proof mechanism, and its outcome is the unique element of the strong core. Under weak preferences, the literature has proposed several generalizations of TTC to preserve Pareto efficiency and strategy-proofness. The procedures of these algorithms are similar in that their each step essentially consists of three stages. In the departure stage, a group of agents are removed with their assignments, if there are no beneficial trades involving them in subsequent steps. In the pointing stage, a rule selects a unique pointee for each agent such that at least one beneficial trading cycle is formed. These selection rules are crucial for ensuring strategy-proofness. In the trading stage, cycles formed in the pointing stage are traded.
Our definition of GTTC is similar to those procedures, but it does not specify a selection rule for the pointing stage. It only requires that at least one beneficial trading cycle is formed . We prove that this requirement is sufficient to ensure that the outcome of the algorithm belongs to the rectified exclusion core. If someone wants to achieve strategy-proofness, a specific selection rule in the existing algorithms can be employed.
Generalized Top Trading Cycle
@@
  • Pointing: Let each agent point to one of the others who hold his favorite objects such that at least one beneficial trading cycle is generated; that is, at least one agent in the cycle strictly prefers the object held by the pointee to the object he holds.
  • Trading: Clear the cycles generated in the Pointing stage such that every agent in every cycle obtains the object held by the pointee. Go to the next step.
In each step of GTTC, the pointing stage generates at least one beneficial trading cycle. By trading the cycle, at least one agent obtains a strictly better object than the object he holds. Since there are finite agents and finite objects, GTTC must stop in finite steps. When GTTC stops, all agents are involved in departing groups. These groups can be ordered as a sequence $(G_{1},G_{2},\ldots,G_{k})$ such that each group $G_{\ell}$ departs before the group $G_{\ell+1}$. An agent may become satisfied at some point in the algorithm but departs in a later step. Once an agent becomes satisfied, he remains satisfied in subsequent steps, until he departs with his assignment.
@@
GTTC may not find all elements of the rectified exclusion core. In Example 3, $\delta$ and $\eta$ are in the rectified exclusion core, but they cannot be found by GTTC. [^8] Specifically, in step one of GTTC, $1$ and $3$ must point to $2$ and $2$ must point to one of them. Thus, either $\{1,2\}$ or $\{2,3\}$ form a cycle and exchange their endowments. After that, there is no more beneficial trading cycle. So, GTTC finds either $\mu$ or $\sigma$.
@@ 5 Rectified strong core
Similar to our motivation for the rectified exclusion core, this section proposes a modification of the strong core to address its emptiness issue. Section 4 has shown that BK’s exclusion right provides a useful conceptual tool for explaining coalition formation under the strong core, based on which we propose the condition $\forall i\in C_{\sigma\sim\mu}$, $\mathcal{I}_{i}(\mu(i))\subseteq\omega(C)$” to regulate the blocking behavior of unaffected agents under weak preferences. However, exclusion blocking does not need to satisfy the feasibility condition required by standard core concepts, where a blocking coalition must redistribute their endowments among themselves (i.e., $\sigma(C)=\omega(C)$”). By replacing condition (2) of Definition 5 with this feasibility condition, we obtain the following concept.
@@
Compared to the definition of the strong core, the above definition adds the condition (3). We formulate their difference as follows: the strong core assumes that an unaffected agent is willing to join a blocking coalition if and only if the coalition owns one object that makes him unaffected and assigns the object to him, whereas the rectified strong core assumes that an unaffected agent is willing to join a blocking coalition if and only if the coalition owns all the objects that make him unaffected and assigns one of them to him. Thus, the strong core is a subset of the rectified strong core.
Under strict preferences, this difference vanishes, and the rectified strong core collapses to the strong core. As discussed, $C_{\sigma\sim\mu}$ can be viewed as being compelled to join the coalition due to the threat posed by $C_{\sigma>\mu}$. Condition (3) of Definition 6 generalizes this interpretation to settings with weak preferences, which has been discussed in Section 4. In contrast, the conventional definition of the strong core literally applies the same condition on unaffected agents from strict preferences to weak preferences.
We summarize our results about the rectified strong core in the following proposition.
Proposition 3.
@@
Theorem 1 and Proposition 3 together imply the following relationships between the various core concepts discussed in the paper:
@@
The set inclusion relation between any two core concepts can be strict. In $^{\prime}_{I}$ of Example 1, the strong core is a strict subset of the exclusion core. In Example 3, the nonempty exclusion core is a strict subset of the rectified exclusion core, which equals $\{\mu,\sigma,\delta,\eta\}$. In Example 6 below, the rectified exclusion core is a strict subset of the rectified strong core. In I of Example 1, the rectified strong core, coinciding with the strong core, is a strict subset of the weak core.
@@
We show that $\mu$ is in the rectified strong core, but not in the rectified exclusion core. In $\mu$, the two pairs $\{1,2\}$ and $\{3,4\}$ respectively exchange their endowments.
@@
To show that $\mu$ is in the rectified strong core, suppose that $\mu$ is rectification blocked by a coalition $C$ via another $\mu^{\prime}$. Since $\mu$ is Pareto efficient, it must be that $C\subsetneq I$. Since $2$ is indifferent between all objects, it does not hold that $\mathcal{I}_{2}(\mu(2))\subseteq\omega(C)$. Therefore, $2\notin C$. Since $1$ and $3$ are the only agents who have not received their most preferred objects in $\mu$, either $1\in C_{\mu^{\prime}>\mu}$ or $3\in C_{\mu^{\prime}>\mu}$ (or both). However, because $2\notin C$, if $3\in C$, it is impossible for $3$ to receive $b$ in $\mu^{\prime}$. Therefore, it must be that $1\in C_{\mu^{\prime}>\mu}$. Then, it must be that $\mu^{\prime}(1)=c$, which requires $3\in C$ and $4\notin C$, since $4$ must be worse off in $\mu^{\prime}$. Therefore, $C=\{1,3\}$. But this means that $\mu^{\prime}(3)=a$, which is a contradiction, since $3$ is worse off in $\mu^{\prime}$.
@@
We then show that $\mu$ is rectification exclusion blocked by $C=\{1,3\}$ via $\sigma$. Note that both agents in the coalition are strictly better off in $\sigma$, and $4$ is the only agent who is worse off in $\sigma$. It is easy to verify that $\mu(4)=c\in\{a,c\}={}^{*}(\{1,3\}|\omega,\mu)$.
6 Housing market with multiple copies of objects
This section examines a special case of the housing market model in which agents’ indifferent preferences stem from the existence of multiple copies of objects. Formally, we consider a setting in which there exists a finite set of object types, denoted by $\mathcal{O}$. For each $x\in\mathcal{O}$, let $O_{x}\subseteq O$ denote the set of the copies of $x$. Therefore, $O=\cup_{x\in\mathcal{O}}O_{x}$. Each $i\in I$ owns an object $\omega(i)\in O$ and has a preference relation i over $O$ such that, for any distinct $o,o^{\prime}\in O$, $o\sim_{i}o^{\prime}$ if and only if $o,o^{\prime}\in O_{x}$ for some $x\in\mathcal{O}$; that is, $o$ and $o^{\prime}$ are different copies of the same object type. Let $I_{x}$ denote the set of owners of the copies of $x$.
In this special model, the strong core may still be empty. In Example 3, $a$ and $c$ can be viewed as two copies of the same object type. The strong core is empty in the example.
However, we prove that the exclusion core is nonempty. Notably, the exclusion core coincides with the set of outcomes of TTC in an artificial priority-based allocation model.
Specifically, let each $x\in\mathcal{O}$ rank the agents in $I_{x}$ using a priority order x, where $i{}_{x}j$ means that $i$ is ranked above $j$. Given a priority structure $({}_{x})_{x\in\mathcal{O}}$, TTC proceeds as follows: in each step, among the remaining agents and object types, let each agent point to his most preferred object type, and each object type point to its highest-priority owner; in each generated cycle, let each agent receive the object owned by the highest-priority owner of the object type he points to, and then remove them.
This algorithm is a refinement of the generalization of TTC introduced by abdulkadirouglu2003school for the school choice model. Their algorithm does not distinguish between the different copies of the same object type, whereas to obtain our result, we need to specify the copy of the object type each agent receives. By varying the priority structure, we obtain different outcomes of TTC. We prove that an allocation belongs to the exclusion core if and only if it can be found by TTC under some priority structure.
Proposition 4.
In the housing market model with multiple copies of objects, the exclusion core equals the set of the outcomes of TTC in the artificial priority-based allocation model.
We illustrate this proposition by revisiting Example 3.
Example 3 revisited (Finding the exclusion core).
We view $a$ and $c$ as two copies of an object type $x$, and view $b$ as the unique copy of another object type $y$.
In the first step of our TTC, $1$ and $3$ point to $y$, and $2$ points to $x$. If $1$ is ranked above $3$ in the priority order for $x$, then $1$ and $2$ form a cycle, with $1$ receiving $b$ and $2$ receiving $a$. This finds the allocation $\sigma$ in the exclusion core. If, instead, $3$ is ranked above $1$, we then find the other allocation $\mu$ in the exclusion core.
However, if we assign arbitrary copies of object types to agents involved in cycles, we then also find $\delta$ and $\eta$, which are not in the exclusion core but are equivalent to its elements.
In this special model, the exclusion core is still not equivalence-closed (see Example 3). Since the rectified exclusion core is equivalence-closed, it must include all allocations that are equivalent to the elements of the exclusion core. These allocations are exactly those found by abdulkadirouglu2003school’s TTC where agents can be assigned arbitrary copies of the object type they point to when involved in cycles. The set of these allocations can be viewed as the “equivalence closure” of the exclusion core. However, Example 7 below shows that this closure may still be strictly smaller than the rectified exclusion core; that is, the rectified exclusion core may have elements that are not equivalent to any element of the exclusion core.
Example 7 (Closure of exclusion core rectified exclusion core).
Consider the following market with three object types and five agents. Object $a$ is the only copy of an object type. Objects $b$ and $b^{\prime}$ are two distinct copies of another type, and $c$ and $c^{\prime}$ are two distinct copies of the third type.
11 22 $2^{\prime}$ 33 $3^{\prime}$
$\omega$: $a$ $b$ $b^{\prime}$ $c$ $c^{\prime}$
$\mu$: $b$ $c$ $c^{\prime}$ $a$ $b^{\prime}$
$\sigma$: $b$ $a$ $c$ $b^{\prime}$ $c^{\prime}$
1 2 $_{2^{\prime}}$ 3 $_{3^{\prime}}$
$b,b^{\prime}$ $a$ $a$ $a$ $a$
$a$ $^{\prime} c,cc,c^{\prime}$ $^{\prime} b,bb,b^{\prime}$
$c,c^{\prime}$ $b$ $b$ $^{\prime} c,cc,c^{\prime}$
We show that $\mu$ is in the rectified exclusion core, but it is not equivalent to any element of the exclusion core.
Suppose that $\mu$ is rectification exclusion blocked by a coalition $C$. In $\mu$, only one agent in $\{2,2^{\prime},3^{\prime}\}$ can be made strictly better off, and the agent must receive object $a$ to be better off. Thus, to block $\mu$, $3$ must be evicted from $a$. If $1\in C$, since he must be unaffected, all owners of his indifferent objects must belong to $C$. So, $\{2,2^{\prime}\}\subseteq C$. Since at least one of $2$ and $2^{\prime}$ must be unaffected, all owners of his indifferent objects must belong to $C$. So, $\{3,3^{\prime}\}\subseteq C$, which, however, is impossible. Thus, $1\notin C$, which implies that $C$ must indirectly control $a$. So, $C$ must directly or indirectly control both $b$ and $b^{\prime}$. If $C$ directly controls $b$ and $b^{\prime}$, it must be that $\{2,2^{\prime}\}\subseteq C$. Then, as above, it must be that $\{3,3^{\prime}\}\subseteq C$, which is impossible. If $C$ indirectly controls $b$ and $b^{\prime}$, it must be that $\{3,3^{\prime}\}\subseteq C$, which again is impossible.
In TTC, under any priority structure, in the first step, $1$ points to the object type of $b$ and $b^{\prime}$, and both $2$ and $2^{\prime}$ point to $a$. Thus, one of $2$ and $2^{\prime}$ must receive $a$, depending on whose priority is higher. For instance, if $2$ is ranked above $2^{\prime}$ in the priority order, we obtain the allocaiton $\sigma$. However, $\mu$, in which neither of $2$ and $2^{\prime}$ receives $a$, cannot be generated by TTC and is not equivalent to any element of the exclusion core.
7 Concluding remark
@@
This section concludes the paper by examining other potential solution concepts for the housing market model. The first solution we consider is the vNM stable set. A set of allocations is a vNM stable set if each allocation inside the set is not dominated by any other allocation inside the set (internal stability), and each allocation outside the set is dominated by an allocation inside the set (external stability). wako1991some shows that the strong core, when it is nonempty, is the unique vNM stable set based on weak domination. However, wako2007nonexistence shows that, in general, a vNM stable set based on either strong or weak domination does not exist. Several papers modify the definition of weak domination and then prove that the corresponding vNM stable set coincides with the set of competitive allocations. However, as discussions in the related literature section, the set of competitive allocations has undesirable features.
@@
demuynck2019myopic introduce the myopic stable set (MSS) for a general class of social environments. A MSS always exists and is unique for a social environment with a finite state space. In the housing market model, a weak domination MSS is a set of allocations satisfying three conditions: (1) no coalition of agents can benefit from deviating from an allocation inside the set to an allocation outside the set (deterrence of external deviations); (2) from any allocation outside the set there is a finite sequence of coalition blocking that leads to an allocation inside the set (iterated external stability); (3) no strict subset of the set satisfies the former two conditions (minimality). demuynck2019myopic have proved that, in the housing market model under strict preferences, the weak domination MSS coincides with the strong core. However, we show that, when preferences are weak, the weak domination MSS may include Pareto inefficient allocations.
Example 2 revisited (MSS includes Pareto inefficient elements).
@@
The weak domination MSS = $\{\mu,\mu^{\prime},\sigma,\sigma^{\prime}\}$. So, it has Pareto inefficient elements, $\mu^{\prime}$ and $\sigma^{\prime}$. To verify that this set is a MSS, note that $\mu$ is weakly dominated by $\mu^{\prime}$, $\mu^{\prime}$ is weakly dominated by $\sigma$ and $\sigma^{\prime}$, $\sigma$ is weakly dominated by $\sigma^{\prime}$, $\sigma^{\prime}$ is weakly dominated by $\mu$ and $\mu^{\prime}$, $\delta$ is weakly dominated by $\mu$ and $\sigma$, and $\omega$ is weakly dominated by any other allocation.
Note that both the vNM stable set and the MSS are setwise solutions; whether an allocation belongs to the set depends on other allocations in the set. In contrast, the various core concepts discussed in our paper are pointwise solutions; whether an allocation belongs to a core is a property of the allocation, irrespective of other allocations in the core.
yilmaz2022stability propose a solution based on the bargaining set developed by aumann1964bargaining. Their idea is that, when a coalition blocks an allocation to reach a new allocation, it should consider the possible counter-blocking of the new allocation by other coalitions. An allocation is in the bargaining set if whenever it is blocked by coalition, there exists a counter-blocking of the new allocation by another coalition that overlaps with the original coalition, and the new coalition blocks the new allocation by claiming their welfare in the original allocation. When a coalition $C$ weakly blocks an allocation $\mu$ via another $\sigma$, there is a degree of freedom to select $\sigma$. yilmaz2022stability impose the selection rule such that, in $\sigma$, all agents outside $C$ who are affected by the blocking receive their own endowments, and the others who are unaffected keep their assignments in $\mu$. The bargaining set is nonempty and Pareto efficient, and lies between the strong core and the weak core. However, it may be strictly larger than the strong core when the latter is nonempty (Example 8). The bargaining set and our solutions do not include each other.
@@
The strong core = $\{\delta\}$. We show that $\mu$, which is not in the strong core, is in the bargaining set. In $\mu$, $2$ receives his endowment, and the remaining agents exchange endowments along a cycle $1\rightarrow 5\rightarrow 4\rightarrow 3\rightarrow 1$. There exists only one coalition $\{1,2\}$ that weakly blocks $\mu$ via an allocation $\sigma$ in which $1$ and $2$ exchange endowments. By the selection rule of yilmaz2022stability, the other agents receive their own endowments in $\sigma$. Then, $\sigma$ is weakly blocked by $\{2,4,5\}$ via an allocation $\mu^{\prime}$ in which $2$ receives his endowment, and $4$ and $5$ exchange their endowments. In $\mu^{\prime}$, all members of $\{2,4,5\}$ receive objects indifferent to their assignments in $\mu$, and the counter-blocking coalition includes one member of the initial blocking coalition (i.e., $2$). Therefore, $\mu$ is in the bargaining set defined by yilmaz2022stability.
@@
A Proofs of Propositions 1, 2, 3, and 4
(If) Suppose that $\mu$ is a Pareto efficient allocation, and there exists a partition of agents $T=(T_{1},T_{2},\ldots,T_{t})$ such that, for every $T_{k}\in T$, $\mu(T_{k})=\omega(T_{k})$, and, for every $i\in T_{k}$, $\mu(i)\in B_{i}\big(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})\big)$. We prove that $\mu$ is in the exclusion core.
Suppose that $\mu$ is exclusion blocked by some coalition $C$ via some $\sigma$. Since $\mu$ is Pareto efficient, there must exist an agent who becomes worse off in $\sigma$ compared to $\mu$. Thus, $C$ must evict some agent. However, we prove that this is impossible. First, all agents in $T_{1}$ cannot join $C$, because all of them have obtained their favorite objects in $\mu$ and thus cannot be made strictly better off. Given this, because all agents in $T_{2}$ have obtained their favorite objects among $O\backslash\omega(T_{1})$, if any agent in $T_{2}$ joins $C$, the agent must become strictly better off by obtaining an object owned by some agent in $T_{1}$. However, because $\mu(T_{1})=\omega(T_{1})$, $C$ does not control $\omega(T_{1})$ and thus cannot evict any agent in $T_{1}$. So, $T_{2}$ cannot join $C$. This argument can be inductively applied to all remaining groups in $T$ to conclude that no agents can join $C$. Therefore, $\mu$ is not exclusion blocked.
(Only if) Suppose that the exclusion core is nonempty, and $\mu$ is any allocation in it. So, $\mu$ is Pareto efficient. Moreover, in $\mu$, agents can be partitioned into disjoint groups such that the agents in each group trade their endowments along a cycle represented by

### v3

@@ 4.1 An algorithm to find elements of the rectified exclusion core
We present an algorithm called generalized top trading cycles (GTTC) and prove that all allocations it finds belong to the rectified exclusion core.
Under strict preferences, TTC is an individually rational, Pareto efficient, and strategy-proof mechanism, and its outcome is the unique element of the strong core. Under weak preferences, the literature has proposed several generalizations of TTC to preserve the three properties. The procedures of these algorithms are similar in that their each step essentially consists of three stages. In the departure stage, a group of agents are removed with their assignments, if there are no beneficial trades involving them in subsequent steps. In the pointing stage, a rule selects a unique pointee for each agent such that at least one beneficial trading cycle is formed. These selection rules are crucial for ensuring strategy-proofness. In the trading stage, cycles formed in the pointing stage are traded.
Our definition of GTTC is similar to those algorithms, but it does not specify a selection rule for the pointing stage. It only requires that at least one beneficial trading cycle is formed. We prove that this requirement is sufficient to ensure that the outcome of the algorithm belongs to the rectified exclusion core. If someone wants to achieve strategy-proofness, a specific selection rule in the existing algorithms can be employed.
Generalized Top Trading Cycles
@@
  • Pointing: Let each agent point to the owner of one of his favorite objects among the remaining ones, such that at least one beneficial trading cycle is generated; that is, some agent in the cycle strictly prefers his pointee’s object to the object he holds.
  • Trading: Clear the beneficial trading cycles generated in the Pointing stage such that every agent in every cycle obtains his pointee’s object. Go to the next step.
In each step of GTTC, the pointing stage generates at least one beneficial trading cycle. By trading the cycle, at least one agent obtains a strictly better object than the object he holds. Since there are finite agents and finite objects, GTTC must stop in finite steps. When GTTC stops, all agents are involved in departing groups. These groups can be ordered as a sequence $(G_{1},G_{2},\ldots,G_{K})$ such that each group $G_{\ell}$ departs before the group $G_{\ell+1}$. An agent may become satisfied at some point in the algorithm but departs in a later step. Once an agent becomes satisfied, he remains satisfied in subsequent steps, until he departs with his assignment. By varying agents’ pointees and the cycles generated in the pointing stage, GTTC may find different allocations. We prove that all outcomes of GTTC are in the rectified exclusion core.
@@
GTTC may not find all elements of the rectified exclusion core. In Example 3, $\delta$ and $\eta$ are in the rectified exclusion core, but they cannot be found by GTTC. [^9] Specifically, in step one of GTTC, $1$ and $3$ must point to $2$ and $2$ must point to one of them. Then, either $\{1,2\}$ or $\{2,3\}$ form a cycle to exchange their endowments. After that, there is no more beneficial trading cycle. So, GTTC finds either $\mu$ or $\sigma$.
@@ 5 Rectified strong core
Similar to our proposal of the rectified exclusion core, this section proposes a modification of the strong core to address its emptiness issue. Section 4 has shown that BK’s exclusion right provides a useful conceptual tool for explaining coalition formation under the strong core, based on which we propose the condition $\forall i\in C_{\sigma\sim\mu}$, $\mathcal{I}_{i}(\mu(i))\subseteq\omega(C)$” to regulate the blocking behavior of unaffected agents under weak preferences. However, blocking coalitions under the exclusion core may not satisfy the feasibility condition required by standard core concepts that blocking coalitions must reallocate their endowments among themselves. By imposing this feasibility condition while retaining our previous condition on unaffected agents in a blocking coalition, we obtain the following modification of the strong core.
@@
Compared to the definition of the strong core, the above definition adds the condition (3). We formulate their difference as follows: the strong core assumes that an unaffected agent is willing to join a blocking coalition if and only if the coalition assigns an indifferent object to him, whereas the rectified strong core requires that the coalition must own all the objects that make him unaffected and assign one of them to him. So, the rectified strong core is a superset of the strong core. Under strict preferences, this difference vanishes and the rectified strong core collapses to the strong core. As discussed, under strict preferences, $C_{\sigma\sim\mu}$ can be viewed as being compelled to join the coalition due to the threat posed by $C_{\sigma>\mu}$. Condition (3) of Definition 5 generalizes this interpretation from strict preferences to weak preferences. In contrast, the conventional definition literally applies the same condition from strict preferences to weak preferences.
The following theorem is our second main result in the paper.
Theorem 2.
@@
The two theorems together imply the following relationships between the various core concepts discussed in the paper:
@@
The set inclusion relation between any two core concepts can be strict. In $^{\prime}_{I}$ of Example 1, the strong core is a strict subset of the exclusion core. In Example 3, the nonempty exclusion core is a strict subset of the rectified exclusion core. In the following Example 6, the rectified exclusion core is a strict subset of the rectified strong core. In I of Example 1, the rectified strong core, coinciding with the strong core, is a strict subset of the weak core. Other than the weak core, the remaining core concepts coincide with the strong core when it is nonempty. In particular, they coincide under strict preferences.
@@
We show that $\mu$ is in the rectified strong core but not in the rectified exclusion core. In $\mu$, the two pairs $\{1,2\}$ and $\{3,4\}$ respectively exchange their endowments.
@@
To show that $\mu$ is in the rectified strong core, suppose that $\mu$ is rectification blocked by a coalition $C$ via another $\mu^{\prime}$. Since $\mu$ is Pareto efficient, it must be that $C\subsetneq I$. Since $2$ is indifferent between all objects, $2\notin C$. Since $1$ and $3$ are the only agents who can be made strictly better off, either $1\in C_{\mu^{\prime}>\mu}$ or $3\in C_{\mu^{\prime}>\mu}$ (or both). However, if $\in C, 33$ must receive $b$ in $\mu^{\prime}$, which contradicts $2\notin C$. So, it must be that $1\in C_{\mu^{\prime}>\mu}$. Then, it must be that $\mu^{\prime}(1)=c$, which requires that $3\in C$ and $4\notin C$, since $\mu(4)=c$ and $4$ most prefers $c$. So, $C=\{1,3\}$, and $3\in C_{\mu^{\prime}\sim\mu}$. However, this means that $\mu^{\prime}(3)=\mu(3)=d$, which contradicts $4\notin C$.
@@
We then show that $\mu$ is rectification exclusion blocked by $C=\{1,3\}$ via $\sigma$. Note that both agents in the coalition are strictly better off in $\sigma$, and $4$ is the only agent who is worse off in $\sigma$. It is easy to verify that $\mu(4)=c\in\{a,c,d\}={}^{*}(\{1,3\}|\omega,\mu)$.
6 Other potential solutions
This section examines other potential solution concepts for the housing market model.
The vNM stable set
@@ 6 Other potential solutions
The first potential solution is the vNM stable set. A set of allocations is a vNM stable set if each allocation inside the set is not dominated by any other allocation inside the set (internal stability), and each allocation outside the set is dominated by an allocation inside the set (external stability). Wako (1991) shows that the strong core, when it is nonempty, is the unique vNM stable set based on weak domination. However, Wako et al. (2007) shows that, in general, a vNM stable set based on either strong or weak domination does not exist. Several papers have modified the definition of weak domination and proved that the corresponding vNM stable set coincides with the set of competitive allocations; see discussions in Related literature in Section 1.
Myopic stable set
@@
Demuynck et al. (2019) introduce the myopic stable set (MSS) for a general class of social environments. A MSS always exists and is unique for a social environment with finite states. In the housing market model, a weak domination MSS is a set of allocations satisfying three conditions: (1) no coalition of agents can benefit from deviating from an allocation inside the set to an allocation outside the set (deterrence of external deviations); (2) from any allocation outside the set there is a finite sequence of coalition blocking that leads to an allocation inside the set (iterated external stability); (3) no strict subset of the set satisfies the former two conditions (minimality). Demuynck et al. (2019) have proved that, in the housing market model under strict preferences, the weak domination MSS coincides with the strong core. However, we show that under weak preferences, the weak domination MSS may contain Pareto inefficient allocations.
Example 2 revisited (MSS contains Pareto inefficient elements).
@@
The weak domination MSS = $\{\mu,\mu^{\prime},\sigma,\sigma^{\prime}\}$. So, it has Pareto inefficient elements $\mu^{\prime}$ and $\sigma^{\prime}$. To verify that this set is a MSS, note that $\mu$ is weakly dominated by $\mu^{\prime}$, $\mu^{\prime}$ is weakly dominated by $\sigma$ and by $\sigma^{\prime}$, $\sigma$ is weakly dominated by $\sigma^{\prime}$, $\sigma^{\prime}$ is weakly dominated by $\mu$ and by $\mu^{\prime}$, $\delta$ is weakly dominated by $\mu$ and by $\sigma$, and $\omega$ is weakly dominated by any other allocation.
Note that both the vNM stable set and the MSS are setwise solutions; whether an allocation belongs to the set depends on other allocations in the set. In contrast, the various core concepts we discuss are pointwise solutions; whether an allocation belongs to a core is a property of the allocation, irrespective of other allocations in the core.
Bargaining set
Yılmaz and Yılmaz (2022) propose a solution based on the bargaining set developed by Aumann and Maschler (1964). Their idea is that, when a coalition weakly blocks an allocation to reach a new allocation, we should consider the possible counter-blocking of the new allocation by other coalitions. In a market, an allocation $\mu$ is said to be in the bargaining set if it is in the weak core and whenever it is blocked by a coalition $C$ via another $\sigma$, there exists a coalition $C^{\prime}$ that weakly blocks $\sigma$ via some $\mu^{\prime}$ such that $C^{\prime}\cap C\neq\emptyset$ and for all $i\in C^{\prime}$, $\mu^{\prime}(i)\sim_{i}\mu(i)$. That is, the counter-blocking coalition overlaps with the original coalition and blocks the new allocation by claiming their welfare in the original allocation. When a coalition $C$ weakly blocks an allocation $\mu$ via some new $\sigma$, since there is a degree of freedom to select $\sigma$, Yılmaz and Yılmaz impose the selection rule such that, in $\sigma$, all agents outside $C$ who are affected by the blocking receive their own endowments and the others who are unaffected keep their assignments in $\mu$.
The bargaining set is nonempty and Pareto efficient and lies between the strong core and the weak core. However, Example 7 shows that the bargaining set may be strictly larger than the strong core when the latter is nonempty. So, the bargaining set may have elements not belonging to our solutions. There also exist examples in which some elements of our solutions are not in the bargaining set. [^10] In Example 6, $\mu$ is in the rectified strong core, but we show that it is not in the bargaining set. To see this, note that $\mu$ is only weakly blocked by $\{1,2,3\}$ via $\sigma$, and $\sigma$ is only weakly blocked by $\{2,3,4\}$ via another $\mu^{\prime}$ in which the three agents exchange their endowments in the cycle $3\rightarrow 2\rightarrow 4\rightarrow 3$. However, $\{2,3,4\}$ does not weakly block $\sigma$ by claiming their welfare under $\mu$, since $\mu^{\prime}(3)\succ_{3}\mu(3)$. So, the counter-blocking does not hold and $\mu$ is not in the bargaining set. We can similarly construct an example in which an element of the rectified exclusion core is not in the bargaining set. Another difference between the bargaining set and core concepts is that the bargaining set requires a two-step verification. That is, for every possible coalition blocking, we need to verify the existence of a possible counter-blocking, which might be computationally more challenging than verifying a core concept.
@@
The strong core = $\{\delta\}$. We show that $\mu$, which is not in the strong core, is in the bargaining set. In $\mu$, $2$ receives his endowment, and the remaining agents exchange endowments along a cycle $1\rightarrow 5\rightarrow 4\rightarrow 3\rightarrow 1$. There exists only one coalition $\{1,2\}$ that can weakly block $\mu$ via another $\sigma$ in which $1$ and $2$ exchange endowments. By the selection rule of Yılmaz and Yılmaz (2022), the other agents receive their own endowments under $\sigma$. Then, $\sigma$ is weakly blocked by $\{2,4,5\}$ via an allocation $\mu^{\prime}$ in which $2$ receives his endowment, and $4$ and $5$ exchange their endowments. The counter-blocking coalition $\{2,4,5\}$ overlaps with the initial blocking coalition $\{1,2\}$, and under $\mu^{\prime}$, all members of $\{2,4,5\}$ receive objects indifferent to their assignments under $\mu$. So, $\mu$ is in the bargaining set defined by Yılmaz and Yılmaz (2022).
@@
A Proofs of Lemma 1, Proposition 1, and Proposition 2
(“Only if”) Suppose that the exclusion core is nonempty and $\mu$ is its element. So, $\mu$ is Pareto efficient. In $\mu$, agents can be partitioned into disjoint groups such that the agents in each group trade their endowments in a cycle represented by